The first term of a geometric series is 9 and the ratio of the sum of the first eight terms to the sum of the first four terms is 97:81.




Calculate the first 3 terms of the sequence it is given that all the terms are positive.

1
Expert's answer
2022-01-31T17:03:49-0500

The sum of first n terms of a geometric progression with first term "a" is,"~S_{n} = \dfrac{a(r^{n}-1)}{r - 1}, r\ne1."

Here given

"a = 9, S_{8}:S_{4}=97:81".

Therefore,

"\dfrac{S_{8}}{S_4} = \dfrac{\frac{9(r^{8}-1)}{r - 1}}{\frac{9(r^{4}-1)}{r - 1}}= \dfrac{9(r^{8}-1)}{r - 1}\cdot \dfrac{r - 1}{9(r^{4}-1)}\\
\dfrac{97}{81}=\dfrac{r^{8}-1}{r^{4}-1}\\
\dfrac{97}{81}=\dfrac{(r^{4})^2-1}{r^{4}-1}\\ \dfrac{97}{81}=\dfrac{(r^{4}-1)(r^{4}+1)}{r^{4}-1}=r^4 +1\\
\dfrac{97}{81}-1=r^4\\
\dfrac{16}{81}=r^4\\
(\frac{2}{3})^4=r^4\\
\therefore r=\frac{2}{3}"


Hence the first three terms are "a, ar, ar^2" , that is 9,6,4.